OpenStudy (anonymous):

What is the surface area of a cone that has a radius of 8 feet and a slant height of 12 feet? (Use 3.14 for π.) ?

OpenStudy (jim766):

SA = πr^2 + π r l , l = slant height, r = radis

OpenStudy (jim766):

\[SA = 3.14(8^2)+3.14(8)(12)\]

OpenStudy (jim766):

you got this?

OpenStudy (anonymous):

ya....